### v1

# Coarse-to-Fine Sparse Meta-Embedding

Abstract
The state-of-the-art recommendation systems have shifted the attention to efficient recommendation, e.g., on-device recommendation, under memory constraints. To this end, the existing methods either focused on the lightweight embeddings for both users and items, or involved on-device systems enjoying the compact embeddings to enhance reusability and reduces space complexity. However, they focus solely on the coarse granularity of embedding, while overlook the fine-grained semantic nuances, to adversarially downgrade the efficacy of meta-embeddings in capturing the intricate relationship over both user and item, consequently resulting into the suboptimal recommendations. In this paper, we aim to study how the meta-embedding can efficiently learn varied grained semantics, together with how the fine-grained meta-embedding can strengthen the representation of coarse-grained meta-embedding. To answer these questions, we develop a novel graph neural networks (GNNs) based recommender where each user and item serves as the node, linked directly to coarse-grained virtual nodes and indirectly to fine-grained virtual nodes, ensuring different grained semantic learning, while disclosing: 1) In contrast to coarse-grained semantics, fine-grained semantics are well captured through sparse meta-embeddings, which adaptively 2) balance the embedding uniqueness and memory constraint. Additionally, the initialization method come up upon SparsePCA, along with a soft thresholding activation function to render the sparseness of the meta-embeddings. We propose a weight bridging update strategy that focuses on matching each coarse-grained meta-embedding with several fine-grained meta-embeddings based on the users/items’ semantics. Extensive experiments substantiate our method’s superiority over existing baselines. Our code is available at https://github.com/htyjers/C2F-MetaEmbed.
keywords:
Lightweight Meta-Embedding, Coarse-to-Fine Learning, ID-Based Recommendations
1 Introduction
Recommender systems (RSs) [1, 2, 3] have revolutionized the way we navigate the overwhelming volume of available information and products by tailoring recommendations to individual user preferences. Benefiting from the collaborative filtering, RSs map users and items into high-dimensional latent spaces [4, 5, 6] for personalized recommendations via the strategies such as the dot product [7], multi-layer perceptrons [8], and graph neural networks [9]. However,with the data scale growing up, complexity and resource demands of these systems heavily grows up. This issue is particularly highlighted in ID-based recommendation systems, where entities are used to uniformly represent all users and items, with each requiring a unique embedding vector. Consequently, traditional RSs suffer from substantial memory overhead, making them difficult to scale up. To address the challenge of optimizing memory usage without compromising recommendation accuracy, there has been a shift towards developing lightweight recommender systems [10, 11, 12, 13, 14] via deep neural network [15, 16].
One strategy is to construct the compact on-device session-based recommender systems [43, 17, 18, 19, 20], trained in the cloud and deployed on devices like smartphones. The primary challenge is to reduce the model size to fit resource-constrained devices. To tackle this, [21] proposed to learn the elastic item embeddings, while adapting to device-specific memory constraints without retraining. Additionally, [22] proposed a personalized elastic embedding learning framework to on-device recommendations, by considering user and device heterogeneity. Moreover, [23] suggested semi-tensor decomposition and self-supervised knowledge distillation [24, 25, 26] techniques for constructing ultra-compact on-device recommendation models. However, these on-device session-based recommenders face the daunting task of pinpointing the optimal embedding sizes and ensuring representational accuracy amidst an immense search space. Simultaneously, under memory constraints, learning from a teacher model necessitates denser embeddings, resulting in a higher likelihood of duplicate embeddings. Conversely, opting for a higher quantity of embeddings to reduce redundancy may lead to sparse embeddings, hindering sufficient knowledge to be captured from the teacher model.
To address this issue, parameter-sharing techniques have been proposed, such as hashing-based methods. The basic idea is to represent entities (such as user/item IDs) using a combination of meta-embedding vectors determined by hash functions, which can be shared across entities. By sharing parameters, hashing-based methods [27, 28] significantly reduce the memory footprint while maintaining distinct representations for each entity. These solutions utilize one or several meta-embedding tables, also known as codebooks, comprising fixed-size embedding vectors. Each user/item ID is hashed into a combination of indexes, enabling the construction of entity embeddings by aggregating meta-embedding vectors drawn from these hashed indexes. This approach, known as compositional embedding, allows the meta-embedding tables to contain far fewer embedding vectors than a full embedding table, yet still offer the unique representations to each entity. [29] proposed clustered embedding learning, where entities within the same cluster jointly learn and share an embedding, significantly reducing memory costs. [30] proposed representing each entity by combining a pair of sparse embeddings from two independent, substantially smaller meta-embedding tables, which are encouraged to encode mutually complementary information. [31] introduced a lightweight embedding framework dedicated to graph neural networks (GNNs) based recommenders, which can adaptively learn all meta-embeddings based on semantics and provide weighted meta-embedding assignments for each entity. However, existing meta-embedding methods tend to focus solely on coarse-grained learning, and fail to capture richer, fine-grained semantic nuances. This limitation significantly impacts the efficacy of meta-embeddings in representing the intricate information of the user and item, consequently resulting into the suboptimal recommendations. The above standing-out problem naturally promotes us to delve into the following questions how the meta-embedding can learn different grained semantics and how the fine meta-embedding can strengthen the representation of coarse-grained meta-embedding for ID-based recommendation.
As motivated above, we propose a novel GNN-based model to facilitate multi-grained semantic learning for meta-embedding. Specifically, we construct a graph based on direct interactions between users and items. Following that, we design a two-tiered virtual node structure [44]. First, coarse meta-embedding nodes are superimposed, directly connecting to all real user and item nodes, thus capturing broad, shared semantics. Subsequently, fine-grained nodes are established, which exclusively link to these coarse-grained nodes, enabling the mapping of more detailed semantics. This hierarchical and selective connectivity ensures that the model learns wide-ranging semantic information at the coarse-grained level, while focusing on fine-grained personalized features. This coarse-to-fine learning strategy fosters a well understanding of the semantics of users/items from global and local side.
To effectively combine the two phases of our embedding learning approach, we adopt a process of optimizing the coarse meta-embeddings as a springboard for developing sparse fine-grained embeddings. Through SparsePCA(Sparse Principal Component Analysis) [42], we selectively generate random entries within the coarse-grained meta-embedding codebook. This targeted initialization not only preserves the sparsity but also ensures the associative relationships between users and items. Moreover, we design the soft thresholding during training then dynamically adjusts sparsity levels. Unlike SparsePCA’s rigid approach, soft thresholding adapts to data and enhancing the robustness. This combination establishes a finely-tuned, dynamic embedding process that remains sensitive to fine-grained semantics. Last but not least, a novel weight comes up to bridge update strategy that dynamically matches coarse-grained meta-embeddings with multiple fine-grained meta-embeddings based on semantic relevance. This innovation ensures that our method remains efficient and effective in representing complex information. Extensive experiments demonstrate that our approach outperforms the existing methods in both recommendation accuracy and memory efficiency, making it highly suitable for applications under memory constraints.The overall pipeline is shown in Fig.1.
In summary, our contributions are summarized below:
  • $\bullet$
    We propose a novel GNN-based model for multi-grained semantics of meta-embeddings, the hierarchical and selective connectivity ensures that the model learns wide-ranging semantic information at the coarse level, while also focusing on granular, personalized features at the fine level.
  • $\bullet$
    An initialization approach is adopted by optimizing the coarse meta-embeddings and SparsePCA to selectively seed the embedding codebook, which preserves the sparsity and maintains associative relationships between users and items.
  • $\bullet$
    A soft thresholding technique is developed to dynamically adjust the sparsity levels during training, while offering a more adaptable and robust model that responds to the varied intricacies of the data.
  • $\bullet$
    Lastly, we develop an innovative weight to bridge update strategy that dynamically aligns coarse-grained meta-embeddings with multiple fine-grained meta-embeddings based on semantic relevance, ensuring efficacy of the recommendation.
Figure 1: Illustration of the proposed pipeline. Our basic idea is to construct a the hierarchical and selective connectivity graph (a), which ensures that the model learns wide-ranging semantic information at the coarse level (b), while also focusing on granular, personalized features at the fine level (c)
2 Preliminaries
Before shedding light on our technique, we elaborate the lightweight embedding for ID-based recommendation and the GNN-based meta-embedding learning.
2.1 Efficient Meta-Embedding for ID-Based Recommendations
In ID-based recommendation, with the user set $\mathcal{U}$ and the item set $\mathcal{I}$, the total number of entities is given by $N=|\mathcal{U}|+|\mathcal{I}|$. The user-item interactions are denoted by $\textbf{R}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$, where 0 signifies no interaction, and 1 indicates the presence of an interaction between a user-item pair. To enhance the memory efficiency of the embedding layer, a common approach involves utilizing a compact meta-embedding codebook, i.e., $\textbf{E}_{meta}\in\mathbb{R}^{m\times d}$, with $d$ representing the dimensionality of the embeddings. This codebook serves as a replacement for the more extensive full embedding matrix $\textbf{E}\in\mathbb{R}^{N\times d}$, where the count of meta-embeddings $m$ is significantly smaller than $N$. To generate an embedding for each entity, a common approach involves using an assignment matrix $\textbf{S}\in\mathbb{R}^{N\times m}_{\geq 0}$. Each row of this matrix corresponds to a user or item entity, where the non-zero elements specify which meta-embeddings are selected from the $m$ available options, along with their respective weights in forming the entity’s embedding. Consequently, the compositional embeddings for all entities, denoted as $\widehat{\textbf{E}}\in\mathbb{R}^{N\times d}$, can be calculated as follows:
$\begin{split}\widehat{\textbf{E}}=\textbf{S}\textbf{E}_{meta}.\end{split}$ (1)
Based on the above, recent developments have demonstrated that Graph Neural Networks (GNNs) can effectively propagate collaborative signals among interconnected entities. Leveraging this, GNN-based recommenders construct a user-item interaction graph to encapsulate the semantic correlations present among entities, as evidenced in their interactions.
2.2 Graph-Propagated Meta-Embeddings
The GNN-based recommenders typically adopt an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$ to represent those connections in the user-item interaction graph, facilitating the propagation of collaborative signals among connected entities for embedding learning. The adjacency matrix A is structured as follows:
$\textbf{A}=\begin{bmatrix}0&\textbf{R}\\ \textbf{R}^{T}&0\end{bmatrix},$ (2)
where R represents the user-item interaction matrix. For each user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, their embeddings $\textbf{e}_{u}$ and $\textbf{e}_{i}$ are retrieved from $\widehat{\textbf{E}}$ and serve as inputs to the user-item interaction graph A. To enhance adaptability, a compositional embedding approach that considers semantic relationships among entities and provides a more refined, weighted combination of assigned meta-embeddings for each entity is proposed. As described in [31], meta-embeddings are represented as additional virtual nodes connected to a set of nodes within the user-item interaction graph A. The assignment matrix S represents the connections between entities and the meta-embedding codebook. Integrating S into A results in an expanded interaction graph $\textbf{A}^{\prime}$:
$\textbf{A}^{\prime}=\begin{bmatrix}\textbf{A}&\textbf{S}\\ \textbf{S}^{T}&0\end{bmatrix},$ (3)
where $\textbf{A}^{\prime}\in\mathbb{R}^{(N+m)\times(N+m)}$. The input embeddings (layer-0 embeddings) to the GNN are generated by stacking the compositional embedding $\widehat{\textbf{E}}$ and the meta-embedding codebook $\textbf{E}_{meta}$:
$\textbf{H}_{(0)}=\begin{bmatrix}\widehat{\textbf{E}}\\ \textbf{E}_{meta}\end{bmatrix},$ (4)
with $\textbf{H}_{(0)}\in\mathbb{R}^{(N+m)\times d}$. The propagation operation at layer $l+1$ is defined as:
$\textbf{H}_{(l+1)}=(\textbf{D}^{-\frac{1}{2}}\textbf{A}^{\prime}\textbf{D}^{-\frac{1}{2}})\textbf{H}_{(l)},$ (5)
where $\textbf{A}^{\prime}$ is the adjacency matrix of the expanded interaction graph and $\textbf{D}\in\mathbb{R}^{(N+m)\times(N+m)}$ is the diagonal degree matrix of $\textbf{A}^{\prime}$. $\textbf{D}^{-\frac{1}{2}}\textbf{A}^{\prime}\textbf{D}^{-\frac{1}{2}}$ creates a symmetrically normalized adjacency matrix. The final GNN embeddings are obtained by averaging the embeddings from all layers:
$\textbf{H}=\frac{1}{L+1}\sum_{l=0}^{L}\textbf{H}_{(l)},$ (6)
allowing the graph-propagated representations of the full entity set $\textbf{H}_{full}\in\mathbb{R}^{N\times d}$ and the meta-embedding codebook $\textbf{H}_{meta}\in\mathbb{R}^{m\times d}$ to be retrieved by splitting H as follows:
$\textbf{H}_{full}\leftarrow\textbf{H}[:N,:],\quad\textbf{H}_{meta}\leftarrow\textbf{H}[N:,:].$ (7)
The meta-embeddings $\textbf{H}_{meta}$ incorporate collaborative information from semantically similar user/item neighbors, enabling a more effective meta-embedding assignment for each user/item. For the recommendation task, the user’s graph-propagated embedding $\textbf{h}_{u}$ and the item’s graph-propagated embedding $\textbf{h}_{i}$ are retrieved from $\textbf{H}_{full}$, and their recommendation affinity prediction is calculated as:
$\widehat{y}_{ui}=(\textbf{h}_{u})^{T}\textbf{h}_{i}.$ (8)
This prediction supports the Bayesian personalized ranking loss (BPR) [32], which is used to optimize each entity’s associated meta-embeddings:
$\mathcal{L}_{\textit{BPR}}=\sum_{(u,i^{+},i^{-})\in\mathcal{B}}-\ln{\sigma(\widehat{y}_{ui^{+}}-\widehat{y}_{ui^{-}})}+\lambda||\Theta||^{2},$ (9)
where $\mathcal{B}$ refers to either the entire training set or a training batch, $(u,i^{+},i^{-})$ is a triplet consisting of the sampled user $u$’s observed interacted item $i^{+}$ and unvisited item $i^{-}$, $||\Theta||^{2}$ represents the $L_{2}$ regularization term applied to the trainable parameters, and $\lambda$ determines the regularization weight in the loss function. With a fixed assignment $\mathbf{S}$ in each iteration, the expanded interaction graph $\mathbf{A}^{\prime}$ utilizes the Eq.9 to back-propagation to update the values in the meta-embedding codebook $\mathbf{E}_{meta}$.
2.3 Learning Assignment Weights
As the meta-embedding codebook $\textbf{E}_{meta}$ is learned, it is essential to continually update the weights in S to achieve a more accurate mapping between entities and their meta-embedding assignments. For instance, two closely related entities should have their embeddings composed of a similar set of meta-embeddings, reflecting the core principle of collaborative filtering. However, jointly updating the S and $\textbf{E}_{meta}$ will introduce difficulties in finding a stable solution. To solve this challenges, [31] proposes to preserved the mapping from each entity to its corresponding meta-embeddings after $L$ layers of propagation in the expanded interaction graph $\textbf{A}^{\prime}$ as same as the in the input layer of the GNN. Thus:
$\begin{split}\textbf{H}_{full}=\textbf{S}\textbf{H}_{meta},\end{split}$ (10)